# GraphPro: Graph Pre-training and Prompt Learning for Recommendation

### 20 Dec 2023 v2

@@ 4.2.3. Comparison with Full-Data Training
We place the results and discussion in Appendix A.1.2 due to space limitation.
@@ 4.3.1. Key Components in Fine-Tuning
We conduct a comprehensive ablation study to examine the effectiveness of the key components in the design of GraphPro, both in the pre-training and fine-tuning stages. To facilitate comparisons with the original design, we create three variants of GraphPro, with each variant removing one key component. Specifically, these variants are:
  • (-)PT: We disable the Prompt Tuning module, which utilizes prompt edges derived from historical interactions. Instead, we directly fine-tune the pretrained weights using new edges.
  • (-)GT: We suppress the adaptive GaTing mechanism during the fine-tuning, which endows dynamic knowledge transformation.
  • (-)IU: We exclude the Interplotive Update module and keep the pretrained weights unchanged for each fine-tuning step.
Based on the results in Figure 3, we make the following observations: 1) All three key components contribute positively to our GraphPro design. Removing any component leads to a significant decrease in recommendation accuracy and, in some cases, longer convergence epochs. This demonstrates the effectiveness of these components. 2) The structural prompt and interplotive update mechanisms enhance accuracy and mitigate local optima effects in dynamic learning. Disabling these components allows faster convergence but results in significantly worse accuracy. 3) The adaptive gating mechanism accelerates model convergence by facilitating better gradient discovery, leading to improved accuracy. Removing the gating mechanism results in substantially longer convergence epochs and worse accuracy, suggesting its importance in addressing the distribution gap between the fine-tuning on snapshots.
@@ 4.3.2. Effect of Pre-trained Model
We examine the impact of different pretrained model designs on downstream fine-tuning performance by comparing four variants. These variants investigate the effects of relative time encoding and representation power of the pretrained models. The original model, ”LGN(+)TE,” is trained on LightGCN with time encoding (TE). ”LGN(-)TE” removes the time encoding design during pretraining. Additionally, ”MixGCF(+)TE” and ”SimGCL(+)TE” utilize stronger models for pretraining while keeping the fine-tuning model as LightGCN. Figure 4 presents the performance comparison of both pretraining and fine-tuning stages.
  • The temporal prompt mechanism significantly accelerates convergence and improves prediction accuracy in both pretraining and fine-tuning stages. It effectively guides the model to leverage important temporal information during message passing.
  • Stronger pretrained models yield better performance, aligning with findings in pre-trained language (Devlin et al., 2018) and vision models (Dosovitskiy et al., 2020). This demonstrates the scalability and adaptability of our framework, enabling powerful pretrained models to achieve superior performance. It highlights the potential of leveraging large pretrained models in recommendation tasks, fine-tuning powerful embeddings with lightweight models for downstream benefits.
@@ 4.4. Learning Impact Analysis (RQ5 & RQ6)
This section analyzes how the fine-tuning design benefits the learning of node representations for recommendation. We categorize users into two groups based on whether they undergo fine-tuning during each time step. The untuned users represent cold-start users in that period. On the Amazon data, we evaluate the two user groups separately at each time step, visualizing the results in Figure 5.
  • GraphPro effectively learns refined representations for both tuned and untuned (cold-start) users. It achieves dominant performance for both groups in most cases. This advantage is attributed to the structural prompt design, where previous interactions provide informative knowledge for improving representation learning.
  • GraphPro exhibits greater performance enhancement from long-term tuning and prediction compared to the baseline. While the baseline initially outperform GraphPro in the first snapshot, GraphPro consistently outperforms the strongest baseline from $T_{2}$ to $T_{8}$. This demonstrates the superiority of our GraphPro design in mitigating local optima effects and achieving better long-term gains in dynamic learning scenarios.
@@ 4.4.2. Efficiency in Learning
This section focuses on the learning efficiency of our GraphPro design. GraphPro is a parameter-efficient method that minimizes the number of learnable weights for pretraining and fine-tuning, in contrast to incremental training methods. This leads to minimal additional training cost in terms of time and computation. To demonstrate the efficiency gains and faster convergence of our design, we compare the training curves of GraphPro with the second baseline methods (EvolveGCN-O and GPF) on the Taobao and Koubei datasets, as shown in Figure 6. The results clearly indicate that GraphPro achieves significantly better performance while requiring fewer learning epochs. For instance, GraphPro converges after four fine-tuning stages, consuming approximately half the epochs and training time compared to the baselines. These findings highlight the substantial learning efficiency improvements offered by GraphPro.
@@
Figure 6. The training curves for GraphPro and the baselines on the Taobao and Koubei datasets are shown. The scatters indicate the evolving performances across different stages (pre-training and fine-tuning) with respect to training epochs. The star marker represents the final convergence point, and the right y-axis represents the overall time consumption.
4.5. Online Deployment and A/B Test (RQ7)
We deploy GraphPro on a large-scale online content consumption platforms (specific name withheld due to anonymous requirements) with millions of users, to evaluate its effectiveness in personalized content recommendation. We integrate GraphPro with the main CTR prediction model, utilizing user embeddings trained with unsupervised deep graph infomax (DGI) as pretrained weights. Prompt edges are created using historical item-to-user interactions, and the pretrained user embeddings are fine-tuned with a 1-layer GNN to derive item representations for training the main model. The pretrained embeddings are updated synchronously with the main model at a ten-minute granularity. In the online A/B test, we allocate an equal user engagement of around 2 million users to GraphPro and the online model separately. We evaluate the performance on 4 metrics related to CTR and click count over a 5-day period, as shown in Table 3. The results demonstrate that GraphPro significantly improves the real-world recommender system by effectively modeling evolving user and item representations and leveraging deep user interests through pretraining and fine-tuning. Notably, GraphPro is easy to deploy with minimal effort, making it a cost-effective solution for enhancing online recommender systems.
@@ 5. Related Works
GNNs for Recommendation. GNNs have gained prominence in recommenders for extracting multi-hop collaborative signals (Gao et al., 2023). NGCF (Wang et al., 2019) and PinSage (Ying et al., 2018) are popular GNN models that refine user and item embeddings recursively using message-passing. GCCF (Chen et al., 2020b) introduces a residual structure, while LightGCN (He et al., 2020) simplifies the architecture by removing non-linear transformations. Researchers have also explored extending GNNs to model complex collaborative relationships, using techniques such as hypergraph learning (Yu et al., 2021; Li et al., 2022) and intent disentanglement (Wu et al., 2022; Wang et al., 2022). Self-supervised learning (SSL) has recently been applied to address sparsity issues in graph-based recommenders, with methods like contrastive learning on user-item graphs (Li et al., 2023; Shuai et al., 2022).
@@
Pre-training and Fine-tuning on GNNs. Inspired by successful pre-training and fine-tuning in NLP, researchers have explored empowering GNNs with similar techniques. Strategies like contrastive learning and infomax-based pre-training have been developed for better representation learning (Hu et al., 2019; You et al., 2020; Veličković et al., 2018; Sun et al., 2019a). Pre-training methods, such as link prediction and feature generation, and prompt-based fine-tuning have also been proposed (Hu et al., 2020; Hou et al., 2022; Fang et al., 2022; Liu et al., 2023b; Sun et al., 2023, 2022). However, these approaches have not fully addressed the challenges of pre-training and fine-tuning in dynamic graph learning, leading to suboptimal performance in temporal-transfer tasks. In contrast, our proposed GraphPro excels in temporally dynamic graph learning under the pre-training and fine-tuning paradigm.
For recommendation tasks, (Hao et al., 2021; Liu et al., 2023a) propose pre-training models specifically for user and item modeling. In (Hao et al., 2021), a GNN is pre-trained as a pretext task to simulate the cold-start scenario. In (Liu et al., 2023a), a side-information-based pre-training scheme is designed. However, these methods primarily focus on stationary recommendation scenarios and overlook the time-evolving nature of user preferences. As a result, their generalization as time-aware recommenders is limited.
@@
Dynamic Graph Learning. Learning on temporally evolving graphs is an emerging research trend. Existing approaches, including EvolveGCN (Pareja et al., 2020), Dyngraph2vec (Goyal et al., 2020), DGNN (Manessi et al., 2020), ROLAND (You et al., 2022), and WinGNN (Zhu et al., 2023), employ various techniques such as RNNs, dense layers, and recurrent layers to capture graph dynamics. However, these methods lack a pre-training and fine-tuning framework for dynamic graph learning, and may introduce noisy perturbations to user and item representations. DGCN (Li et al., 2020) incorporates dynamics in graph-based recommendation learning, but it does not explicitly consider a dynamic graph setting with snapshots and is limited to evaluation within a single graph.
Sequential Recommendation. Sequential recommendation, also known as next-item recommendation, is another research area that focuses on temporal-aware recommendation settings. Representative works in this field include i) attention-based methods: SASRec (Kang and McAuley, 2018), BERT4Rec (Sun et al., 2019b), STOSA (Fan et al., 2022); ii) GNN-based recommenders: SURGE (Chang et al., 2021) amd Retagnn (Hsu and Li, 2021); iii) SSL-enhanced models: S3-rec (Zhou et al., 2020) and ICL (Chen et al., 2022). While our approach also considers temporal dynamics, it differs significantly from next-item recommendation methods. Firstly, sequential recommenders typically employ auto-regressive encoders, which limit their capability to accurately predict the next item and are not directly comparable to graph-based methods that can recall top-K items. Secondly, existing sequential methods do not explicitly consider time in terms of daily or weekly intervals. Instead, they retrieve a fixed-length historical sequence.
@@ 6. Conclusion
This study proposes a novel framework, GraphPro, that integrates dynamic graph pre-training with prompt learning to improve the adaptation and scalability of time-aware recommender systems. By employing a temporal prompt mechanism, our framework enables the transfer of valuable knowledge from the pre-trained model to downstream recommendation tasks on newly arrived data. In addition, the inclusion of graph-structured prompt learning with adaptive gating mechanism, allowing for the incorporation of crucial contextual information, facilitating fine-tuning and adaptation to changing behavior dynamics. Through a comprehensive set of experiments on diverse real-world datasets, we demonstrate that GraphPro outperforms state-of-the-art baselines in making dynamic recommendations across different time slots. Our future work lies in investigating the interpretability of the prompt graph edges in GraphPro, which can provide insights into the contextual information that is used to fine-tune the pre-trained GNNs.
@@ Appendix A Appendix
(a) Taobao & Koubei
(b) Amazon
Figure 7. We conducted a comparison between our GraphPro and a full-data training framework using LightGCN. The comparison focused on the average training time for a single epoch, with the results depicted by a horizontal line on the right Y-axis.
A.1. Additional Experiments
A.1.1. Hyper-parameter Sensitivity
Here we study how sensitive GraphPro is towards hyper-parameter settings change. On the three datasets, we include all the hyper-parameters of our GraphPro design, which are time interval $\tau$ for the temporal prompt, updating window $\omega$ for interplotive update, and sampling decay $\phi$ for the structural prompt. The search space varying from datasets are:
@@ A.1.1. Hyper-parameter Sensitivity
We present the results in Figure 8 and summarize our findings as: 1) Generally, GraphPro is more sensitive to these hyper-parameters on Amazon than on Taobao and Koubei. This may due to the longer fine-tuning and predicting span on Amazon (9 weeks). This suggests that the hyper-parameters in GraphPro have larger effect on long-term performances. To obtain better long-term performances, the hyper-parameters should be chosen more carefully. 2) GraphPro is more sensitive to updating window $\omega$ than other HPs. A shorter window size may be insufficient to include useful updates of user and item representations, while a longer one may risk bringing more noise that is relevant to current representation learning. 3) Overall, small adjustments to these hyper-parameters do not lead to significant performance degradation, indicating the robustness of our model to hyper-parameter variations. We recommend selecting suitable hyper-parameters based on the data characteristics.
@@
A.1.2. Comparsion with Full-Data Training
In this section, we investigate whether GraphPro can perform on par or even outperforms the vanilla full-data training method (”FULL”). On all three datasets, we analyze the performances of GraphPro and FULL on every snapshot in terms of Recall and average epoch time. We plot the results in Figure 7, where we can have following observations: 1) On Taobao and Koubei, GraphPro consistently outperforms FULL on all testing snapshots by a significant margin, which is a counter-intuitive finding. This confirms that on these datasets, taking time as an important factor and dynamically learning from temporal snapshots derive better recommendation accuracy. On Amazon, GraphPro and FULL generally perform on par. Specifically, GraphPro wins on the starting snapshots and for the rest, the performances are tightly close, which demonstrates that switching to the dynamic paradigm of GraphPro would yield potential performance improvements with minimal to no noticeable performance degradation, comparing to full training. 2) Our dynamic GraphPro framework greatly reduces the average epoch training time by a great margin, compared to FULL. Concretely the efficiency is boosted by 60x, 24x and 81x separately on Taobao, Koubei and Amazon. We attribute this advantage to the pre-training and prompt learning paradigm design. This advantage makes our model well-suited for efficient recommendation learning and prediction in real-world scenarios.

### 14 Jan 2024 v3

@@ 4.2.3. Comparison with Full-Data Training
We place the results and discussion in Appendix A.2 due to space limitation.
@@ 4.3.1. Key Components in Fine-Tuning
We perform a detailed ablation study to assess the impact of GraphPro’s crucial elements during pre-training and fine-tuning. For comparative analysis against the original design, we introduce three GraphPro variants, each omitting a specific key component:
  • (-)PT: This variant removes the Prompt Tuning component that leverages prompt edges from past interactions, opting to fine-tune the model directly with new interactions.
  • (-)GT: Adaptive GaTing mechanism, which facilitates dynamic knowledge transformation in the fine-tuning phase, is omitted.
  • (-)IU: The Interplotive Update module is excluded, maintaining static pre-trained weights during fine-tuning.
From the data in Figure 3, we discern that: 1) Each of the three integral elements of GraphPro plays a vital role. Omitting any one of them not only diminishes recommendation precision but may also prolong the time to convergence, underscoring their collective efficacy. 2) The structural prompt and interpolative update features are instrumental in boosting accuracy and avoiding local optima during dynamic learning. Their absence can lead to quicker convergence but at the cost of significantly reduced accuracy. 3) The adaptive gating feature is key to expedited convergence and enhanced accuracy through improved gradient flow. Without it, we observe notably extended convergence times and diminished accuracy, highlighting its role in bridging the distribution shifts encountered during fine-tuning across snapshots.
@@ 4.3.2. Effect of Pre-trained Model
We assess how different pretraining model architectures influence fine-tuning outcomes by examining four variants, focusing on the use of relative time encoding and the intrinsic representational strength of the models. The baseline model ”LGN(+)TE” incorporates time encoding (TE) with LightGCN. The ”LGN(-)TE” variant excludes TE during pretraining. The other two, ”MixGCF(+)TE” and ”SimGCL(+)TE,” employ more robust models for pretraining while maintaining LightGCN for fine-tuning. Performance contrasts between pretraining and fine-tuning for these variants are depicted in Figure 4.
  • The temporal prompt mechanism enhances both the speed of convergence and the accuracy of predictions during pretraining and fine-tuning. It directs the model to capitalize on crucial temporal cues in the message-passing process.
  • The use of more robust pretrained models correlates with improved performance, echoing trends seen in pre-trained language models like BERT (Devlin et al., 2018) and vision models like ViT (Dosovitskiy et al., 2020). Our framework’s scalability and adaptability are evident, as it allows the use of powerful pretrained models to excel in recommendation. This underscores the strategy of refining large pretrained models with more streamlined models downstream for enhanced results.
@@ 4.4. Learning Impact Analysis (RQ5 & RQ6)
This section delves into the advantages of our fine-tuning approach on node representation learning for recommendations. We divide users into two categories according to whether they are subject to fine-tuning at each time step. Users who do not undergo fine-tuning are considered as cold-start users for that period. Using the Amazon dataset, we assess these two groups independently at every time step and present the findings in Figure 5.
  • GraphPro effectively enhances representations for both tuned and cold-start users, consistently outperforming in most scenarios. This success is largely due to the structural prompt design, which leverages past interactions to enrich the representations.
  • GraphPro showcases enhanced performance over the baseline with sustained long-term improvement. After an initial lag in the first snapshot, GraphPro overtakes the top baseline from $T_{2}$ through $T_{8}$, underscoring its effectiveness at navigating local optima for superior long-term results in dynamic settings.
@@ 4.4.2. Efficiency in Learning
This section examines the learning efficiency of our GraphPro. As a parameter-efficient approach, GraphPro minimizes learnable weights during both pretraining and fine-tuning, unlike incremental training methods, thereby reducing training time and computational costs. Our comparisons of training curves, shown in Figure 6, reveal that GraphPro not only outperforms EvolveGCN-O and GPF on the Taobao and Koubei datasets but also achieves this with fewer epochs. For example, GraphPro reaches convergence after just four fine-tuning stages, taking roughly half the epochs and time compared to the baselines, underscoring the efficiency improvements GraphPro provides.
@@
Figure 6. The training curves for GraphPro and baselines on the Taobao and Koubei are presented, with scatters marking performance over various stages (pre-training and fine-tuning) against epochs. A star marker denotes the final convergence, and the right y-axis shows the total time consumed.
4.5. Online Deployment (RQ7)
GraphPro is implemented on a major online content platform (name omitted for anonymity), serving millions of users, to assess its effectiveness in personalizing content recommendations. Integrating GraphPro into the core CTR prediction system, it utilizes unsupervised DGI-trained user embeddings as pretraining weights. Historical item-to-user interactions shape prompt edges, and these embeddings are fine-tuned with a 1-layer GNN to inform item representations for the primary model, with updates occurring every ten minutes. During an online A/B test, GraphPro and the current online model each engage approximately 2 million users. Performance is evaluated over a 5-day period using four metrics related to CTR and click count, as presented in Table 3. Results show that GraphPro significantly enhances the real-world recommender system by effectively modeling evolving user and item representations and leveraging deep user interests through pretraining and fine-tuning, all while being easy and cost-effective to deploy.
@@ 5. Related Works
GNNs for Recommendation. GNNs have become key in recommendation systems for capturing multi-hop collaborative signals (Gao et al., 2023). Models like NGCF (Wang et al., 2019) and PinSage (Ying et al., 2018) recursively enhance user and item embeddings through message-passing. GCCF (Chen et al., 2020b) adds a residual structure, whereas LightGCN (He et al., 2020) streamlines by discarding non-linearities. Further innovations include hypergraph learning (Li et al., 2022; Yang et al., 2022) and intent disentanglement (Wu et al., 2022; Ren et al., 2023b) to parse complex collaborative patterns. To tackle sparsity in graph recommenders, self-supervised techniques such as contrastive (Li et al., 2023; Shuai et al., 2022) and generative approaches (Xia et al., 2023) have been introduced.
@@
Pre-training and Fine-tuning on GNNs. Drawing from NLP’s pre-training and fine-tuning success, researchers have applied similar concepts to GNNs. Techniques like contrastive learning and infomax-based pre-training have been designed for advanced representation learning (Hu et al., 2019; You et al., 2020; Veličković et al., 2018; Sun et al., 2019a). Innovations in pre-training, such as link prediction, feature generation, and prompt-based fine-tuning, have emerged (Hu et al., 2020; Hou et al., 2022; Fang et al., 2022; Liu et al., 2023b; Sun et al., 2023, 2022). Yet, these methods haven’t fully overcome the intricacies of dynamic graph learning, often resulting in less-than-ideal performance in tasks with temporal elements. Our GraphPro stands out for its proficiency in handling dynamic graphs within the pre-training and fine-tuning.
In the realm of recommendation tasks, (Hao et al., 2021; Liu et al., 2023a) suggest pre-training models tailored for user and item modeling. (Hao et al., 2021) pre-trains a GNN to mimic cold-start conditions, while (Liu et al., 2023a) develops a pre-training strategy using side information. Nevertheless, these approaches concentrate on static recommendation settings and fail to account for the dynamic evolution of user preferences, thus limiting their effectiveness as time-sensitive recommenders.
@@
Dynamic Graph Learning. Research on learning with temporally evolving graphs is gaining traction. Techniques such as EvolveGCN (Pareja et al., 2020), Dyngraph2vec (Goyal et al., 2020), DGNN (Manessi et al., 2020), ROLAND (You et al., 2022), and WinGNN (Zhu et al., 2023) leverage RNNs and recurrent layers to grasp graph changes. Yet, they miss a dedicated pre-training and fine-tuning strategy specific to dynamic graphs and can introduce noise into user and item representations. DGCN (Li et al., 2020) does consider dynamics in graph recommendation learning, but it falls short of handling graphs with evolving snapshots, limiting its assessment to a static graph.
Sequential Recommendation. Sequential recommendation, a field attuned to time-sensitive recommendation scenarios, showcases works like attention-based approaches: SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019b), GNN-based methods: SURGE (Chang et al., 2021) and Retagnn (Hsu and Li, 2021), and SSL models such as S3-rec (Zhou et al., 2020), ICL (Chen et al., 2022), and DCRec (Yang et al., 2023b). Our approach, while similarly embracing temporal dynamics, diverges from these next-item predictors. Sequential models, with their auto-regressive encoders, face limitations in next-item prediction and differ from graph-based methods that excel in top-K item retrieval. Furthermore, they don’t incorporate explicit time intervals like days or weeks, but rather rely on fixed historical sequences.
@@ 6. Conclusion
This study introduces GraphPro, a new framework combining dynamic graph pre-training with prompt learning, aiming to substantially enhance the adaptability and scalability of recommender systems that are sensitive to temporal dynamics. GraphPro uses a temporal prompt mechanism to transfer knowledge from pre-trained models to recommendation tasks on fresh data. Its graph-structured prompt learning features an adaptive gating mechanism that includes vital contextual information, easing fine-tuning and responsiveness to behavioral shifts. Through an extensive array of experiments conducted on a range of real-world datasets, we have empirically validated that GraphPro consistently outpaces contemporary state-of-the-art baselines, delivering exceptionally accurate dynamic recommendations across various temporal contexts. Looking ahead, our research will delve into the interpretability aspect of GraphPro, particularly focusing on the prompt graph edges.
@@ Appendix A Appendix
A.1. Hyper-parameter Sensitivity
Our research focuses on investigating the response of GraphPro to hyper-parameter adjustments, using three distinct datasets. We conduct a meticulous analysis of the primary hyperparameters that play a crucial role in the operation of GraphPro. These hyperparameters include the temporal prompt interval $\tau$, the updating window $\omega$, which determines the intervals for model updates, and the sampling decay $\phi$, which influences the emphasis on past interactions. To ensure a comprehensive evaluation, we carefully select ranges of values for each hyperparameter that are tailored to the unique characteristics of each experimental dataset utilized.
@@ A.1. Hyper-parameter Sensitivity
Our analysis of hyper-parameter sensitivity in GraphPro is presented in Figure 8, from which we derive several observations:
1) GraphPro exhibits a greater degree of sensitivity to hyperparameter changes when applied to the Amazon data in comparison to those of Taobao and Koubei. This heightened sensitivity on Amazon is likely due to its considerably longer fine-tuning and prediction period, which spans an entire 9 weeks. This duration is substantial when contrasted with the other datasets and suggests a correlation between the span of fine-tuning and the impact of hyperparameters on the model’s long-term performance. For enhanced long-term accuracy, we need to be careful when picking these settings.
2) When examining the details, it becomes evident that GraphPro is notably more affected by the updating window $\omega$ compared to other hyperparameters in its configuration. The size of the updating window requires careful consideration, as it is crucial to strike a delicate balance. If the window is too narrow, the model may fail to incorporate important updates that contribute to the evolving representations of users and items. On the other hand, if the window is too wide, the model may become burdened with irrelevant noise, thus obscuring the current representation learning process.
3) Moreover, our findings reveal a reassuring aspect of GraphPro’s design: its resilience. Minor modifications in the hyper-parameter settings do not precipitate substantial declines in the model’s performance. We emphasize the significance of customizing hyperparameter choices to align with the intrinsic characteristics and unique aspects of the dataset being utilized.
(a) Taobao & Koubei
(b) Amazon
Figure 7. GraphPro v.s. full-data training using LightGCN. The results with the average training time for a single epoch are being depicted by a horizontal line on the right Y-axis.
@@
A.2. Comparsion with Full-Data Training
In this section, our primary objective is to conduct a comprehensive performance comparison between our GraphPro and the traditional full-data training approach (FULL). Our aim is to determine if GraphPro can match or exceed the performance of FULL. To achieve this, we conduct a meticulous comparison using three distinct datasets, with a keen focus on Recall and average epoch time. The results of our analysis are visually represented in Figure 7. Our findings can be summarized into two key observations:
1) In our study, GraphPro stands out on both the Taobao and Koubei datasets, surpassing FULL’s performance at every testing point. This underscores GraphPro’s robust capability in processing time-sensitive data to yield more precise recommendations. With the Amazon dataset, the two approaches are closely matched, but GraphPro shows a distinct advantage in the initial phases. This suggests that GraphPro’s dynamic method can be immediately beneficial, improving performance from the start without losing ground to FULL as the evaluation progresses over time.
2) Furthermore, a significant aspect of GraphPro’s appeal lies in its exceptional efficiency. When compared to the FULL approach, GraphPro cuts down the average training time dramatically, achieving astonishing efficiency gains. Specifically, we observe an impressive 60-fold increase in speed on the Taobao dataset, a 24-fold increase on the Koubei dataset, and an unparalleled 81-fold increase on the Amazon dataset. These improvements in training efficiency can be attributed to the innovative architecture of GraphPro, which seamlessly integrates pre-training and prompt learning techniques. This strategic integration significantly accelerates the learning process, enabling GraphPro to rapidly assimilate and adeptly apply new data to generate accurate and timely recommendations.